36ft fishing line is cut to where the longer length is 6ft longer than the shorter length, what is the length of the shorter line?
Answer by mary164(1)

We have two pieces of fishing line so lets call their length x and y.
Their sum is equal to 36, so x+y=36.
Let x be the shorter length and y the longer.
If we express the value of y in terms of x we get the equation y=X+6. Because x is the shorter length and y is 6 more than the shorter length.
Now we replace y in our equation and solve for x.
X+y=36
x+(x+6)=36 collect like terms
2x+6=36 minus 6 in both sides
2x=30 divide by 2 in both sides
x=15
so the shorter piece of fishing line is 15 ft long.